OpenStudy (anonymous):

A fitnss club charges members a monthly fee and a fee for each fitness class attended. Last month, one member attended 8 fitness classes and paid $80. Another member attendd 14 fitness classes and paid $98. Determine the monthly fee and the fee per fitness class.

OpenStudy (anonymous):

So there's a monthly fee (let's call it 'F') and there's fixed per-class-fee (how about 'C' ?) 1st mameber attends 8 classes and pays $80 - so: \[8 C + F = $80\] 2nd member attends 14 classes and pays $98 - so: \[14 C + F = $98\] so you got two equations in two unknowns. Think you can solve it on your own?

OpenStudy (anonymous):

no, please help through the rest :P

OpenStudy (anonymous):

Use system of elimination to solve for C. -1*(8C + F) = -1*($80) *Multiply by -1 to eliminate F once you add the (14C + F) = $98 equations together So, you now have: 6C = 18 which means C = 3. *Plug those values into the original equations to solve for F. I hope that helps.

OpenStudy (anonymous):

yes it did actually, thanks a ton! =)